A board that is 12 feet long must be cut into two pieces that have lengths in a ratio of 3 to 2. Find the length of the 2 pieces.
ratio 3:2
3/2 = 12/x answer would be 3x = 24 = x = 8. So the sides would be 8 & 4 but the ratio is 2:1, not 3:2...

Answer by macston(5194)

The ratio of length 3:2 3+2=5 The 3 is 3/5 of the total and the 2 is 2/5 and the total is 12 so 3/5(12)=7.2 feet and 2/5(12)=4.8 feet. 7.2 ft+4.8=12 ft and
7.2:4.8 is 3:2